OpenStudy (anonymous):

The sum of three numbers is 122. The third number is 8 less than the first. The second number is 4 times the third. What are the numbers?

Mehek (mehek14):

x + 4(8-x) + 8-x = 122 this is the equation

OpenStudy (anonymous):

Ok, how do i get x?

OpenStudy (anonymous):

what would be your first step in solving for x?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

if i give you 2x + 4 = 10, could you solve for x?

OpenStudy (anonymous):

x=3?

OpenStudy (anonymous):

thats correct, how did you get that answer?

OpenStudy (anonymous):

btw x = 3 may or may not be what x equals in the other equation

OpenStudy (anonymous):

that is just a side example

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

but how did you get the value of x in my example?

OpenStudy (anonymous):

10-4=6 6/2=3

OpenStudy (anonymous):

ah very nice so you do know how to do this

OpenStudy (anonymous):

ok so do the same thing for the other equation

OpenStudy (anonymous):

isolate x on one side of the equation

OpenStudy (anonymous):

first off i would eliminate all the parenthesis by distributing, can you do that and rewrite it?

OpenStudy (anonymous):

x + 4(8-x) + 8-x = 122, distribute the middle term, and rewrite

OpenStudy (anonymous):

lol yeah thats true, so why were u asking how to find x if you know how? lol anyways nice, now just plug x back in to each term to see what each number is

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

first number = x second number = 4(8-x) third number = 8-x